Foto Company makes 40,000 units per year of a part it uses in the products it manufactures. The cost per unit of this part is shown below:

Direct Materials 12
Direct Labor 10.10
Variable Overhead 5.60
Allocated Fixed Overhead 10.50
Total 38.20

An outside supplier has offered to sell Foto Company 40,000 of these parts for $36.70 per unit. If the company accepts this offer, the facilities now being used to make the part could be used to make more units of a product that is in high demand. The additional contribution margin earned on this other product would be $310,000 per year.

Calculate the selling price per unit charged by the outside supplier that would make Foto economically indifferent between making and buying the part.
Answer #1

Calculate selling price charged by outside supplier at indifferent point

Manufacturing cost = Purchase cost

27.70*40000+310000 = 40000X

Selling price = 1418000/40000 = 35.45